2.) an electron moving through an electric field of 500 V/m and a magnetic field of 0.10T experiences the two fields and the electron's direction of motion are mutually perpendicular what is the speed of the electron ?
1
Expert's answer
2012-11-12T11:44:19-0500
Lorentz force is equal to centrifugal acceleration * mass
mv^2/r = qvB
Hence
r = mv/(qB) = 9.1*10^(-31)*10^7/(1.6*10^(-19)*0.02) = 0.003м